A normal distribution has μ = 80 and σ = 10. What is the probability of randomly selecting X < 85 


1
Expert's answer
2021-06-03T15:20:56-0400

"P(X<85)=P(Z<\frac{85-80}{10})=P(Z<0.5)=0.6915."